Presentation is loading. Please wait.

Presentation is loading. Please wait.

1 Offline status report LNF 19 Sep 07. 2 News:  MC04  all LSF=1 completed (~535 pb -1 )  mrc dst of 2005 with correct   +  - e + e - ~completed.

Similar presentations


Presentation on theme: "1 Offline status report LNF 19 Sep 07. 2 News:  MC04  all LSF=1 completed (~535 pb -1 )  mrc dst of 2005 with correct   +  - e + e - ~completed."— Presentation transcript:

1 1 Offline status report LNF 19 Sep 07

2 2 News:  MC04  all LSF=1 completed (~535 pb -1 )  mrc dst of 2005 with correct   +  - e + e - ~completed (1.13 fb -1 done, 95%).  MC all_phys2 (dtr_gb_nr=1) started for 2004 data: 400 pb done in ~15 days, 26pb -1 /day  New request (P. Beltrame): production of 250 pb -1 Bhabha at 1 GeV,  =420 nb, LSF = 2. Estimate rate 60pb -1 /day, 1 week to complete.  “Organizational” news: R. Versaci deputy convener

3 3 MC activity: summary of the productions done in 2006/07 All_phys LSF=1 (2004/05) Run Categories Luminosity (pb -1 ) dst K ± 1497/1741 (87%) dst K /1741 (99%) dst  1738/1741 (99%) dst rad1738/1741 (99%) Rare channels (2005) Specific MC prod. (dst’s produced) Luminosity (pb -1 ) K ± rare, LSF= /1215 K S rare, LSF= /1215 K L rare, LSF= /1215  ( e + e -  GeV, LSF=6 254/265  ( e + e -  1GeV, LSF=6 265/265 Production of 1.74 fb -1  all 2004/05 (LSF=1) on 136 slots (  280 B80 CPU). Production completed in ~4 months of running, average rate 18 pb -1 /day. The inefficiency for K ± dst’s mostly caused by the 2GB limit for single file. They will be recovered in future. Production of ~1.2 fb -1 (2005 data) of: K ± rare decays (K ±  e, , e ), K S rare decays (K S  e, ,  e, , 3  K L rare decays (K L , ,  e,  e,  Production completed in 8 days of running. Production of 240 pb 1GeV of:  ( e + e -  ),  =50 nb, LSF=6  ( e + e -  ),  =30 nb, LSF=6 Production completed in 5 days of running.

4 4 Status of reconstruction at Sep 07 DBV ScanComment bad wires ! 215 Various significant ECL mods Minor changes to recon No bias sample for rad Stable ECL Old bias sample for rad Stable ECL for  s = m f ufo as bias sample Stable ECL for off-peak ufo as bias sample Stable ECL for off-peak ufo as bias sample analyzed Fully cal trgmon luminosity in pb -1 by year and DBV First 125 pb -1 of 2004 show large time jitter. We will reprocess if needed. OK Don’t have   +  - e + e - selection. We want to reprocess them

5 5 Short term activities:  Duplicate MC04/05  all statistics (~1.7 fb -1 ) ~120 days (started)  Reprocessing of 2005 data with DBV<25: ~80 days (1.1 fb -1 )  Dst’s holes  Rad MC LSF = 5: ~10 days  Beltrame’s request: 250pb -1 of Bhabha at 1 GeV (~1 week) Within this schedule we should be able to finish this activity in ~7 months from now (i.e. March 08)

6 6 Spares

7 7 Program Events (10 9 ) LSF Time (B80 days) Size* (TB)  all 2004/05 data K  rare decays  2005 data K L rare decays  2005 data K S rare decays  2005 data  ( e + e -  GeV  ( e + e -  GeV Monte Carlo production 2006/07 Program Events (10 9 ) Time (B80 days) Size* (TB) Simulation of all 2001/02 data All the Monte Carlo produced before 2006 *w/out including MC dst’s

8 8 Data 2001/02 (TB) RAW101 Recon55 DST data8 Total164 MC Data 2004 (TB) RAW54 Recon35 DST7 Total96 Tape Storage (Sep 07) 217 TB Data 2005/06 (TB) RAW95 Recon96 DST23 Total214 Luminosity Tag=OK (pb -1 ) Scan+OP282 Total2632 Tot ~700 TB MC dst’s (TB) 2001/ Total36

9 9 Offline resources: Disk space DSTs cached on nfs-mounted disks for fast analysis access CurrentFinal 40 TB data 36 TB MC 76 TB total 40 TB data 80 TB MC* 120 TB total DST size Current DST cache capacity: New purchases: 20 TB FC disk available (16.5 effective) Installed other 30TB Sata2 (22 effective) 38.5 TB available for disk cache 11/19 TB in use for AFS storage *all_phys at LSF=2 + specific productions

10 10 Offline resources: CPU NodesTypeCPUs B80 Equival. fibm14-15,17-34P3 375 MHz 80 fibm35-44P GHz 4096 fibm45-47 (Jan ’06)P5 1.5 GHz SMT “96”184 All360 KLOE CPU accounting is performed in B80 units 1 “B80 CPU” = 1 P3 CPU installed in B80 server

11 11 Attivita’ Offline 2006/07 Attivita’ svolta nel 2006: Ricostruzione dati Off-Peak (~280 pb -1 ), fino a Marzo Preparazione e test della simulazione MC dei dati 2004/05 (~1.8 fb -1 ) Simulazione MC  all dei dati 2005 (~1.2 fb -1 ), Agosto-Dicembre Attivita’ svolta nel 2007: Riprocessamento parte dei dati Off-Peak (~60 pb -1 ), Gennaio Chiusura “buchi” della ricostruzione 2005 (~100 pb -1 ) Riprocessamento dati 2004 (~535 pb -1 ) Produzione MC canali specifici (K rari,  ) Simulazione MC  all dei dati 2005 (~535 pb -1 ), Giugno-Luglio Attivita’ da svolgere nel 2007/08: Duplicazione statistica MC  all 2004/05 (~1.7 fb -1 ) Riprocessamento dati 2005 (~1.1 fb -1 ) Chiusura buchi dst’s, produzione MC di canali specifici, altro…


Download ppt "1 Offline status report LNF 19 Sep 07. 2 News:  MC04  all LSF=1 completed (~535 pb -1 )  mrc dst of 2005 with correct   +  - e + e - ~completed."

Similar presentations


Ads by Google